Tommys Tile Service is planning on purchasing new tile cleaning equipment that will improve their ability to remove tough stRequired 1 Required 2 Required 3 Determine the companys current fixed costs. Fixed costRequired 1 Required 2 Required 3 Determine the companys new break-even point in sales. Break-even point in salesRequired 1 Required 2 Required 3 After the purchase of the equipment, how much revenue does the company need to generate a pr
Answer #1

1) fixed cost= break even point* contribution %

= 331,200*25%

= 82,800

2) break even point= fixed cost/ contribution%

= (82800+8000)/25%

= 363,200

3) sales= profit+ fixed cost/ contribution %

= 105,000+(82800+8000)/25%

= 195,800/25%

Sales = 783,200